OpenStudy (anonymous):

Help me Please!!! ' if f(x)=x^1/2-x and g(x)=2x^3-x^1/2-x, find f(x)-g(x)

OpenStudy (lochana):

f(x) - g(x) =x^1/2-x - (2x^3-x^1/2-x) = x^1/2 - x - 2x^3 + x^1/2 + x = -2x^3 + 2x^1/2

OpenStudy (anonymous):

Thank you for the help!!(:

OpenStudy (lochana):

bye